Determine the kenetic energy of an 875 g chicken moving at a speed of 4.023 m/s
1
Expert's answer
2018-03-08T10:16:07-0500
By the definition of the kinetic energy we have:
KE=1/2 mv^2,
here, m is the mass of the chicken, v is the speed of the chicken.
Then, we get:
KE=1/2 mv^2=1/2∙0.875 kg∙(4.023 m/s)^2=7.1 J.